OpenStudy (anonymous):

•In a paragraph, describe how to simplify each expression: 1.161/2

OpenStudy (anonymous):



OpenStudy (anonymous):

1.161= ?

OpenStudy (anonymous):



OpenStudy (anonymous):

In a paragraph, describe how to simplify each expression

OpenStudy (anonymous):

Amm... that wasnt clear but we can write it : 1.161=1161/1000 k ?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So we have : 1161/1000/2=1161/1000*1/2 Right ?:)